OpenStudy (anonymous):

find the mean and model class Class boundaries Frequency 0.5–3.5 11 3.5–6.5 12 6.5–9.5 4 9.5–12.5 2 12.5–15.5 1

OpenStudy (anonymous):

can anyone tell me the equation of how to find the answer? please

OpenStudy (john_es):

You should find first a point that represents the interval. So to each interval exists one point, the mid point, whose form is, \[x_i=(a+b)/2\] Once you get the midpoint for each interval, you can find the mean, with the formula, \[\overline{x}=\frac{x_1\cdot f_1+x_2\cdot f_2+x_3\cdot f_3+\ldots+x_n\cdot f_n}{n}\]

OpenStudy (john_es):

Also the modal class is the interval with highest frequency. In this case, [3.5,6.5].

OpenStudy (anonymous):

I am a little confused. how do I find the class boundary

OpenStudy (anonymous):

the class boundary is already given to me but I have to figure the mean the full question is class boundary frequency 0.5-3.5 11 3.5-6.5 12 6.5-9.5 4 9.5-12.5 2 12.5-15.5 1 I have to find the mean and the modal.

OpenStudy (john_es):

Well, you need the class average to find the mean with the formula I wrote.

OpenStudy (john_es):

For example, the class average for the first interval is, \[(0.5+3.5)/2=2\]

OpenStudy (anonymous):

so I will do that all of those class boundaries and then divide by 5?

OpenStudy (john_es):

You find all the class average, and you'll obtain a new table.

OpenStudy (john_es):

You don't need to divide by 5 the class average.

OpenStudy (anonymous):

i don't understand. I thought to find the mean you

OpenStudy (john_es):

class average frequency 2 11 5 12 8 4 11 2 14 1

OpenStudy (anonymous):

i had to all of the values and then divide

OpenStudy (anonymous):

add

OpenStudy (anonymous):

oh

OpenStudy (john_es):

Now, multiply each class average by its frequency, add, and divide by the sum of all frequencies. \[Mean=\frac{2\cdot 11+5\cdot 12+8\cdot 4+11\cdot 2+14\cdot 1}{30}\]

OpenStudy (anonymous):

ok so where did you come up with 30?

OpenStudy (john_es):

It is the sum of all frequencies.

OpenStudy (anonymous):

I am totally confused now. This is my first class of statistics and I am completely lost.

OpenStudy (john_es):

Take a tour on this webpage, may be it helps you http://www.mathsteacher.com.au/year8/ch17_stat/03_freq/freq.htm#M1

OpenStudy (anonymous):

ok thanks i will do that.

OpenStudy (anonymous):

I appreciate your help

OpenStudy (john_es):

You're welcome. ;)

OpenStudy (anonymous):

btw my answer is 5 right?